If a batter runs 19mph how many seconds does it take to run 90 feet after he hit a ball 245.7 feet
Answer by jim_thompson5910(35256)

The "he hit a ball 245.7 feet" bit of info is irrelevant because we don't know any of the throwing speeds of the outfielders or the basemen. We also don't know anything about how fast the ball is flying out and how long it takes to land. So we'll ignore that part.

So let's just assume he runs 19 mph and he runs 90 ft. How long does it take? Well let's find out.

-------------------------------------------------------------------------

90 ft * (1 mi/5280 ft) = 90/5280 = 0.017045 miles

90 ft = 0.017045 miles approximately

-------------------------------------------------------------------------

d = r*t

0.017045 = 19*t

0.017045/19 = t

0.000897 = t

t = 0.000897

The approximate time in hours is 0.000897. Let's convert to seconds

0.000897 hours * (60 min/1 hr)*(60 sec/1 min) = 0.000897*60*60 = 3.2292

Keep in mind that much of these calculations are approximate.

So it takes roughly 3.2292 seconds for the batter to run 90 feet. This is assuming the batter maintains a speed of 19 mph.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------